Calculate the ratio of effusion rates of C12 to F2
1
Expert's answer
2019-10-07T06:42:50-0400

Note: I assume you meant "C{l_2}" , not "C12" .


From the kinetic theory we know that effusion rate of component "i" is proportional to it's molar mass (it's known as Graham's Law)


"{Q_i} \sim \frac{1}{{\sqrt {{M_i}} }}"

Thus


"\frac{{{Q_{C{l_2}}}}}{{{Q_{{F_2}}}}} = \sqrt {\frac{{{M_{{F_2}}}}}{{{M_{C{l_2}}}}}} = \sqrt {\frac{{{M_F}}}{{{M_{Cl}}}}} = \sqrt {\frac{{18.998[\frac{{\text{g}}}{{{\text{mol}}}}]}}{{35.446[\frac{{\text{g}}}{{{\text{mol}}}}]}}} \approx 0.732"